# EDITOR: Effective and Interpretable Prompt Inversion for Text-to-Image Diffusion Models

### v1

@@ Appendix C Ablation studies
CLIP Similarity $\uparrow$ 0.784 0.796 0.799 0.802
@@
Prompt Perplexity $\downarrow$ 70.364 80.659 83.457 87.907
Impact of the Number of Iteration. We explore the impact of the number of iteration during the optimization for embedding inversion in Appendix C. Generally, increasing the number of iterations improves the similarity between the generated image and the target image, which is the goal of our optimization. However, the improvement slows down over time, meaning that additional iterations result in smaller gains. Additionally, with more iterations, the quality of the inverted prompt decreases. This happens because over-optimization introduces complex or unclear tokens and semantic structures, making the prompt harder to interpret. To balance image similarity, prompt quality, and computational cost, we set the default number of iterations to 1000.
Impact of the Number of Denoising Step. EDITOR can invert the prompts for text-to-image diffusion models with small denoising steps. We explore the impact of denoising step number from 5 to 15 steps. As shown in Appendix C, the difference in performance across different step numbers is relatively small, indicating that the number of denoising steps has limited influence on the image similarity and the quality of prompts. However, using fewer steps, such as 5 steps, significantly reduces optimization complexity and computational cost, improving overall efficiency. In Stable Diffusion, denoising steps below 5 often lead to insufficient noise reduction, which impacts image quality. Therefore, we set the default number of denoising step to 5.
Table 11: Impact of denoising step.
Metric 5 steps 10 steps 15 steps
@@
Table 12: Impact of different image captioning model.
@@
LPIPS Similarity $\downarrow$ 0.427 0.414 0.467
@@
Impact of Different Image Captioning Models. To justify the image captioning model choices for the prompt initialization, we show the results in Appendix C using three state-of-the-art open-source models: GIT-large [43], BLIP-large [48] and BLIP2-OPT-2.7B [42]. Among the models, GIT-large shows the weakest performance. Notably, BLIP2-OPT-2.7b leverages a large language model, resulting in more sophisticated initialized prompts that achieve better textual alignment with the ground-truth. BLIP-large achieves better image quality with higher CLIP Similarity. We select BLIP-large as our default image captioning model. It is more lightweight compared to BLIP2-OPT-2.7b while still maintaining excellent performance.
Impact of Different Image Captioning Models. To justify the image captioning model choices for the prompt initialization, we show the results in Appendix C using three state-of-the-art open-source models: GIT-large [43], BLIP-large [48] and BLIP2-OPT-2.7B [42]. Among the models, GIT-large shows the weakest performance. Notably, BLIP2-OPT-2.7b leverages a large language model, resulting in more sophisticated initialized prompts that achieve better textual alignment with the ground-truth. BLIP-large achieves better image quality with higher CLIP Similarity. We select BLIP-large as our default image captioning model. It is more lightweight compared to BLIP2-OPT-2.7b while still maintaining excellent performance.
Impact of Random Noise Initialization. The choice of random noise seed in diffusion models significantly affects the output. We further investigated the influence of seed values used for initializing random noise on the results of our method. As shown in Table 13, the impact is minimal. This demonstrates the robustness of our approach and the ability of the model to effectively manage and mitigate the effects of random noise during the inversion process.
@@
Appendix D Efficiency
Table 14: Efficiency comparison between PH2P and EDITOR.
PH2P EDITOR
3482s 1369s
To run an inversion with 1000 iterations on 1 A100 GPU, the optimization times are shown in Table 14. EDITOR improves optimization efficiency by more than double compared to PH2P. We did not compare PEZ because it is based on CLIP inversion, whereas PH2P and EDITOR are based on Stable Diffusion.
Appendix E Broader Impacts
As the value of prompts in text-to-image generation models increases,The risk of prompt stealing, which could infringe upon the intellectual property of prompt engineers and undermine the commercial model of the prompt market [6, 10].In this paper, we propose a method for prompt inversion, which can reverse a target image into a high-quality prompt on text-to-image (T2I) generation models. We are confident that our approach can contribute to a better understanding of this emerging issue and provide solutions to protect the intellectual property of prompt engineers.
Our method also shows potential applications in data attribution and model provenance. The intuition is that if an image is generated by a specific diffusion model or originates from its training data, performing prompt inversion on that image using its source model will yield higher image similarity. By designing appropriate validation pipeline, we can use EDITOR to achieve data attribution and model provenance.
Appendix F Limitations
EDITOR produces prompts that are more interpretable and accurate than those of existing methods, and the resulting prompts can generate images more accurately than advanced image captioning models or visual large language models (vLLMs). However, more advanced diffusion models, such as Stable Diffusion 3 [49], utilize both the T5 model and a CLIP-based encoder model to achieve text-controlled generation. The EDITOR method can extend its capabilities to perform embedding optimization under multiple text encoders, enhancing its generality and performance on advanced models.

### v2

@@ Appendix E Ablation studies
CLIP Score $\uparrow$ 0.784 0.796 0.799 0.802
@@
Prompt PPL $\downarrow$ 70.364 80.659 83.457 87.907
Impact of the Number of Iteration. We explore the impact of the number of iteration during the optimization for embedding inversion in Table E. Generally, increasing the number of iterations improves the similarity between the generated image and the target image, which is the goal of our optimization. However, the improvement slows down over time, meaning that additional iterations result in smaller gains. Additionally, with more iterations, the quality of the inverted prompt decreases. This happens because over-optimization introduces complex or unclear tokens and semantic structures, making the prompt harder to interpret. To balance image similarity, prompt quality, and computational cost, we set the default number of iterations to 1000.
Impact of the Number of Denoising Step. EDITOR can invert the prompts for text-to-image diffusion models with small denoising steps. We explore the impact of denoising step number from 5 to 15 steps. As shown in Table E, the difference in performance across different step numbers is relatively small, indicating that the number of denoising steps has limited influence on the image similarity and the quality of prompts. However, using fewer steps, such as 5 steps, significantly reduces optimization complexity and computational cost, improving overall efficiency. In Stable Diffusion, denoising steps below 5 often lead to insufficient noise reduction, which impacts image quality. Therefore, we set the default number of denoising step to 5.
Table 12: Impact of the number of denoising step.
Metric 5 denoising steps 10 denoising steps 15 denoising steps
@@
[…]
Table 13: Impact of image captioning models.
@@
LPIPS Score $\downarrow$ 0.427 0.414 0.467
@@
[…]
Impact of Different Image Captioning Models. To justify the image captioning model choices for the prompt initialization, we show the results in Table E using three state-of-the-art open-source models: GIT-large (Wang et al., 2022), BLIP-large (Li et al., 2022a) and BLIP2-OPT-2.7B (Li et al., 2022b). Among the models, GIT-large shows the weakest performance. Notably, BLIP2-OPT-2.7b leverages a large language model, resulting in more sophisticated initialized prompts that achieve better textual alignment with the ground-truth. BLIP-large achieves better image quality with higher CLIP Similarity. We select BLIP-large as our default image captioning model. It is more lightweight compared to BLIP2-OPT-2.7b while still maintaining excellent performance.
Impact of Different Image Captioning Models. To justify the image captioning model choices for the prompt initialization, we show the results in Table E using three state-of-the-art open-source models: GIT-large (Wang et al., 2022), BLIP-large (Li et al., 2022a) and BLIP2-OPT-2.7B (Li et al., 2022b). Among the models, GIT-large shows the weakest performance. Notably, BLIP2-OPT-2.7b leverages a large language model, resulting in more sophisticated initialized prompts that achieve better textual alignment with the ground-truth. BLIP-large achieves better image quality with higher CLIP Similarity. We select BLIP-large as our default image captioning model. It is more lightweight compared to BLIP2-OPT-2.7b while still maintaining excellent performance.
Impact of Random Noise Initialization. The choice of random noise seed in diffusion models significantly affects the output. We further investigated the influence of seed values used for initializing random noise on the results of our method. As shown in Table E, the impact is minimal. This demonstrates the robustness of our approach and the ability of the model to effectively manage and mitigate the effects of random noise during the inversion process.
Impact of Beam Search Size in E2T) model. We further analyze the effect of the beam search size in the embedding-to-text decoding stage. As shown in Table E, different beam sizes lead to only marginal variations across similarity, alignment, and perplexity metrics. This demonstrates that the beam search setting has a negligible effect on our final results, confirming the stability of our method with respect to decoding choices. We set the default beam search size in the embedding-to-text decoding stage to 4.
@@
[…]
Table 15: Impact of beam search size.
Metric 1 4
[…]
Appendix F The Use of Large Language Models (LLMs)
In preparing this paper, we employed a large language model (LLM) to aid in editing and polishing the writing. The LLM was used to improve clarity, grammar, and academic style, as well as to suggest alternative phrasings for certain sections such as the contributions, ablation study descriptions, and reproducibility statement. The technical content, research ideas, experiments, and analyses were fully developed and conducted by the authors. No LLM was used to generate experimental results, implement algorithms, or design the methodology.